OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 5) and (0, −4)

OpenStudy (michele_laino):

here we have to apply this formula: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 4 - 5}}{{0 - 2}} = ...?\]

OpenStudy (anonymous):

-9/-2?

OpenStudy (michele_laino):

being \((x_1,y_1)=(2,5)\) and \((x_2,y_2)=(0,-4)\)

OpenStudy (michele_laino):

yes! correct the requested slope, is \(m=9/2\)

OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A. 2 B.-4/5 C.5/4 D. -2

OpenStudy (michele_laino):

here, again we have to apply the same formula above, so we have this: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{1 - \left( { - 5} \right)}}{{4 - 1}} = ...?\]

OpenStudy (anonymous):

6/3

OpenStudy (michele_laino):

correct! \(m=2\)

OpenStudy (anonymous):

Thank You

OpenStudy (michele_laino):

:)